OpenStudy (anonymous):

GIVING MEDALS!!!!!! Find the surface area of a pyramid with a square base with dimensions 100 mm by 100 mm and height of 75 mm.

OpenStudy (anonymous):

erm... are there options?

OpenStudy (anonymous):

No, but i got 28027.76 ?

OpenStudy (anonymous):

I don't know how to do this sorry :( @sourwing @jdoe0001 @MathLegend help?

OpenStudy (jdoe0001):

do you know what a square pyramid looks like?

OpenStudy (anonymous):

I couldn't tell if height meant slant height or other

OpenStudy (anonymous):

Height is from the point of the triangle to the middle of the square on the bottom. Slant height is one of the 4 lines connecting to the top to a corner on the square.

OpenStudy (jdoe0001):

you're asked to find the "surface area" notice the pyramid has 4 triangular faces and 1 squarish bottom the surface area is, the SUM of all those areas the bottom is simple, is a square, and you're given the dimensions the triangular faces, you're given the "base" of the triangle but notice the picture, you can just use pythagorean theorem to get the "slant height" of the pyramid, which is really the "height" of the triangular face once you have the "height" of the triangular face, recall that area of a triangle = \(\bf \frac{1}{2}\cdot base\cdot height\)

OpenStudy (anonymous):

so 4 triangles slant height = sqrt ( 75^2 + 50^2 ) triangle = (1/2) (slant height) (base length) = (1/2) (sqrt ( 75^2 + 50^2 )) (100) base square = (100) (100) so surface area = (4*(1/2)*(sqrt(8125))*(100)) + 1000 = ?

OpenStudy (jdoe0001):

kinda forgot to put the 50 in the picture... but anyhow, just use the pythagorean theorem to find the slant height :)

OpenStudy (anonymous):

@jigglypuff314 do you understand?

jigglypuff314 (jigglypuff314):

they've pointed in the right direction :) "surface area = (4*(1/2)*(sqrt(8125))*(100)) + 1000 = ? "

OpenStudy (anonymous):

19027.76

jigglypuff314 (jigglypuff314):

yes :) that's what I got at least :P

OpenStudy (anonymous):

What about this one? Find the volume of a pyramid with a square base with dimensions 100 mm by 100 mm and height of 75 mm.

OpenStudy (jdoe0001):

\(\bf \textit{volume of a pyramid}=\cfrac{1}{3}\cdot \textit{area of base}\cdot height \)

jigglypuff314 (jigglypuff314):

as in area of base = (100)(100) so volume = (1/3) (100)(100) (75) = ?

OpenStudy (anonymous):

250,000 ?

jigglypuff314 (jigglypuff314):

yeah :)

jigglypuff314 (jigglypuff314):

oh crap, which made me just realize that there was a typo in the other one, sorry surface area = (4*(1/2)*(sqrt(8125))*(100)) + 10000 = ? ^missed a zero

OpenStudy (anonymous):

28027.76 :) And if you look further in the convo... I said that earlier lol

jigglypuff314 (jigglypuff314):

lol XD >,< Ima just a bit cross eyed :P

OpenStudy (anonymous):

Find the radius of the base of a cone with volume equal to 144π in3 and height equal to 12 in

jigglypuff314 (jigglypuff314):

volume = (1/3) (area of base) (height) area of base = pi r^2 volume = (1/3) (pi r^2) (height) 144pi = (1/3) (pi r^2) (12) solve for r

OpenStudy (anonymous):

What is it? I can't seem to find an answer.. ;o

jigglypuff314 (jigglypuff314):

(12/3) pi r^2 = 144pi divide both sides by 4pi r^2 = 36 sqrt both sides r = sqrt 36 = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Last one :) What is the volume of a pyramid that has a base with an area of 81 square feet and a height of 4 feet?

OpenStudy (anonymous):

108 right?

jigglypuff314 (jigglypuff314):

volume of pyramid = (1/3) (area of base) (height) = (1/3) (81) (4) 108 is correct :)

OpenStudy (anonymous):

kk :) ty! Night! :D